Consider the reaction: NaNO3(s) + H2SO4(l) → NaHSO4(s) + HNO3(g) ΔH° = 21.2 kJ How much heat must absorbed by the reaction system to convert 100g of NaNO3 into NaHSO4(s)? 


1
Expert's answer
2021-09-06T02:38:17-0400

molar mass of  NaNO3=85g / mol (23+14+48)

"mass\space of \space NaNO3\space (from\space balanced\space equation)=1 \times 85=85g"

since,15g of NaNO3 is converted =21.2 KJ absolute

"hence, \space100g\space of\space NaNO3\space converted = \frac{100 \times 21.1}{85}=24.9KJ"